OpenStudy (anonymous):

Which equation has both 4 and -4 as possible values of y? A. y3 = 64 B. y2 = 8 C. y3 = 8 D. y2 = 16

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[4^2=16\\ (-4)^2=16\]

OpenStudy (misty1212):

they are the solutions to \(y^2=16\)